OpenStudy (bohotness):

HELP WILL MEDALND FAN HELP WITH 25 QUESTIONS

OpenStudy (bohotness):

The Empire State Building is 1,250 feet tall. The Sears Tower is 204 feet taller than the Empire State Building. Which equation could be used to find h, the height of the Sears Tower? A. 1250 = 204h B. 204 = 1250 + h C. 1250 = 204 + h D. h = 1250 + 204

OpenStudy (bohotness):

i picked c

OpenStudy (bohotness):

@!Call~Me~Addi! @!Call~Me~Addi!